Circular No. 6434

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ak IN NGC 5021 F. Colas, Institut d’Astrophysique de Paris, reports the discovery by

Christian Buil, on CCD images obtained on July 13.93 UT with a 2.12-m reflector at Pic du Midi Observatory, of a supernova (mag 18, compared with Guide Star Catalogue stars) located 25′′ west and 13′′ south of the center of NGC 5021 (α = 13^{h}12^{m}.1, δ = +46^{o}12′, equinox 2000.0). Confirmation of SN 1996ak was reported by M. Reugnier at Toulouse on July 14.8 and by C. Cavadore at Pic du Midi on July 14.9.

A. V. Filippenko, University of California at Berkeley (UCB), reports that his inspection of uncalibrated CCD spectra (range 320–1000 nm) obtained by D. Stern and L. Maxfield (UCB) on July 15 UT with the Lick Observatory 3-m Shane reflector reveals that SN 1996ak is a type-II supernova. The Hα profile consists of a strong, broad emission line and a relatively weak P-Cyg absorption component, but absorption is more prominent among the higher-order Balmer lines.

HV 5756 The MACHO collaboration (cf. IAUC 6312, plus K. Pollard) has dis-

covered that the 17.5-day period type-II Cepheid variable HV 5756 in the Large Magellanic Cloud is an eclipsing binary whose next primary minimum will occur on 1996 July 18.0 UT ± 2 days. The star is located at α = 5^{h}19^{m}26^{s}.8, δ = −69^{o}51′52′′ (equinox 2000.0), with magnitude V ≃ 15.8, V − R ≃ +0.4, and a V semi-amplitude of 0.35 mag. During eclipse, the star will be ∼ 0.5 mag fainter at V. A finder chart and further information about eclipse development may be found at

either http://wwwmacho.mcmaster.ca/HV5756/index.html or http://
wwwmacho.anu.edu.au/HV5756/index.html. Potential observers are re-

quested to contact welch@physics.mcmaster.ca (or Doug Welch at telephone 905-525-9140, x23186) to coordinate observations.

SS LACERTAE P. B. Etzel and N. H. Volgenau, San Diego State University, commu-

nicate: “Pursuant to IAUC 6429, two spectra of SS Lac taken around July 9.37 UT (at historical orbital phase 0.32) provide a velocity difference of 125 km/s, with component strengths reversed from our previous measure. Binarity is confirmed.”

1996 July 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green